OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (0, −4) and a directrix of y = 4. A.f(x) = −16x2 B.f(x) = 16x2 C.f(x) = −one sixteenth x2 D.f(x) = one sixteenthx2

OpenStudy (anonymous):

@random231 @aaronq @sourwing @princeharryyy @satellite73 @DarkMoonZ @Baby_Bear69

OpenStudy (welshfella):

the general form is x^2 = 4ay where a = the y coordinate of the focus can you work it out ?

OpenStudy (anonymous):

what do i work out sorry i have no idea what to do with these types of questions

OpenStudy (welshfella):

the focus is at (0,-4) - that is the y coordinate is - 4

OpenStudy (welshfella):

x^2 = 4(-4)y

OpenStudy (welshfella):

this type of parabola as the origin at (0,0) because x = 0 when y = 0 and as the focus is negative 4 opens downwards

OpenStudy (anonymous):

so x^2=16y?

OpenStudy (welshfella):